OpenStudy (anonymous):

Solve the following system of equations: 7x - y = 4 and 3x + 2y = 9 A. (2, 10) B. (1, 3) C. (5, -3) D. (4, 24) I think its A or D im not sure @iGreen. @Loser66

OpenStudy (anonymous):

https://www.desmos.com/calculator/lyglb7ciw4 The solution is the point where the two lines cross..and where's that?

OpenStudy (anonymous):

@tiffany__ What did you get?

OpenStudy (anonymous):

5&3

OpenStudy (loser66):

@iGreen. I think you should guide her solving the problem in algebraic way. Not all of students know how to solve it by graph. I know it is good but graphing is not the best choice for them.

OpenStudy (anonymous):

Yes, good point.

OpenStudy (loser66):

Here, we use online graphing tool to draw the equations out in 1 second, on test, they have to draw them by hand and it is not easy to not make a mistake which can lead to wrong answer.

OpenStudy (anonymous):

Okay, let's use the substitution method instead. 7x - y = 4 Subtract 7x to both sides: -y = -7x + 4 Multiply -1 to both sides: y = 7x - 4 Now plug in 7x - 4 for y in the 2nd equation: 3x + 2(7x - 4) = 9 Distribute 2 into the parenthesis: 3x + 14x - 8 = 9 Combine like terms: 17x - 8 = 9 Add 8 to both sides, what's 9 + 8?

OpenStudy (anonymous):

@tiffany__

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Yep that gives us: 17x = 17 Now divide 17 / 17, what do you get?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yep, so that's your 'x' term..so which option has a '1' in the beginning?

OpenStudy (anonymous):

@tiffany__

OpenStudy (anonymous):

b @iGreen.

OpenStudy (anonymous):

Yep, so that's your answer. @tiffany__